From acd@roe.ac.uk Thu Dec 15 22:13:15 2016 Date: Thu, 15 Dec 2016 22:13:03 +0000 From: Clive Davenhall To: Bob Mann , Andy Lawrence , "Read, Mike" , Eckhard Sutorius , Mark S. Holliman , Nicholas Cross , Nigel C. Hambly , Rob Blake , Ross Collins , Stelios Voutsinas , Dave Morris , Clive Davenhall Cc: "dante@astro.puc.cl" , James Dunlop , "j.p.emerson@qmul.ac.uk" , ktn keithnoddle.org , "Lorenzo.Rimoldini@unige.ch" , "mcioni@aip.de" , "matt.jarvis@astro.ox.ac.uk" , "mike@ast.cam.ac.uk" , "nigel.metcalfe@durham.ac.uk" , "norman@astro.gla.ac.uk" , "pwl@star.herts.ac.uk" , "rgm@ast.cam.ac.uk" , "Simon.Dye@nottingham.ac.uk" , "s.j.warren@ic.ac.uk" , "tom.shanks@durham.ac.uk" Subject: WFAU VDFS Science Archives progress meeting minutes, 9 December 2016. WFAU VDFS Science Archives progress meeting minutes, 9 December 2016 Present: ETWS, RSC, MSH, MAR, RGM, RPB, NJC, ACD, NCH (part of time) Apologies: STV, DMR, KTN, AL DoNM: 10:00 am Friday 6th January 2017, Villas Meeting Room MAR in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: NJC to complete the deep-stacking software for VVV. Done. The software was completed and working satisfactorily. NJC was recopying from Chile the deep stacks that it was to process. Actions carried forward from 11/11/2016 meeting: --------------------------------------------- ACTION: RGM to investigate the size and content (how does it differ from DR9?) of SDSS DR10 to see if it is worth us hosting and whether it is necessary to keep all of the earlier releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. The part of this action that continues is that RGM had not yet established which early releases could safely be scrapped. RGM also noted that that DR11 and DR12 were now available and now need to be looked at. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan DR13 is installed but not yet cross-mounted. However, we still need to decide which SDSS releases to keep so the present action should remain. ACTION: MAR to write some H-ATLAS documentation once H-ATLAS is available. Continues. ACTION: RPB to check the log files and inform NJC of the problems with the previously listed non-survey releases. Continues. ACTION: NJC to investigate and fix the problems with the previously listed non-survey releases. Continues. This action cannot progress until RPB checks the log files and passes on details of the problems (action above). ACTION: RPB to check with CASU whether they had any WFCAM observations made during the gap in the Edinburgh tapes. It was suggested we might like to check here first! Continues. ACTION: ETWS to ask for and re-transfer the missing SDSS databases. Continues, ETWS now has access to the SDSS directories but cannot find the necessary files. ACTION: MSH to investigate the speed of disk44 Continues. MSH had completed cursory investigations and found no obvious problems. Additional checks are needed. ACTION: ACD to set-up a WIKI page on MS SQL Server on Linux and all to add their input about what and how it needs to be tested for our uses. Continues. Specific points and new actions: -------------------------------- Project management: * RGM noted all WFAU staff would receive their annual PDR before Christmas and that the next WFAU Quarterly Meeting would also be held before the break. ACTION: RGM to complete and circulate the Agenda for the Quarterly Meeting. WFCAM, VISTA and VST updates: * Nothing to report. ACTION: MAR to arrange the next regular CASU/WFAU operations meetings with Mike Irwin. Networking: * The networks were operating nominally. MSH reported that the current switch would remain operational until the proposed new 10 Gbit switch was installed. WSA/VSA/OSA/GES Operations: * ETWS reported: - Calculation and update of the OSA illumination corrections has completed. Also all the objIDs have now been updated. - the OSA VPHAS re-seaming is now 57% complete. - CU3 ingests for the April to June VISTA data have nearly finished. - New browser pages have been completed for the latest database release. * RPB reported that the WSA ingest database was now relocated on ramses16. * ACD noted that a revised version of the documentation for the GES iDR5 linelist had been received and RSC had added it to the Web pages. * MAR added that Maria Rosa had asked that the VMC survey include a list of SMASH / Gaia cross-neighbours. The details were still being discussed. (for SMASH see http://adsabs.harvard.edu/abs/2014AAS...22325444O). Hardware and Systems: * MSH reported that the new NAS boxes had been installed in their racks. RPB inquired whether the boxes could provide the LSST with the approx. 100 Tbyte of temporary staging space needed to hold PANSTARRS data prior to their ingest into a database. MSH replied that sufficient space should be available until next summer. RGM then noted that the PANSTARRS data release was scheduled for the week beginning 12/12/16 but some data may be added later. The data would be retrieved during the spring of 2017 and might arrive on disk rather than being copied across the networks. * RSC was briefly questioned on the reliability of the backups of the SuperCOSMOS archives. He confirmed that both the catalogues and the images were backed-up and current plans envisage making a new backup every year. ACTION: RGM and ACD to discuss the WFAU backup strategy once the current set of backups are complete. Review of Software Development Deadlines: * The list of software deadlines was briefly discussed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/PrepareSoft Other Software: * ETWS reported that he had installed the new GAaP shapelets code on unas for NJC. * NJC noted that he had made some improvements to the matched-aperture photometry and mosaicing software. Review of Survey Release Deadlines: * Progress towards forthcoming survey deadlines was reviewed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Survey Data Release: * The VIKING P96 Team Release had been made on 3/12/16. Non-survey Data Release: * Nothing to report. Astrogrid deployment, VO & Data Analysis services: * Nothing to report. Miscellaneous: * NJC was in discussion with Jim Dunlop about providing details of potential UltraVISTA VSA and ESO releases, in particular about the transfer of OB data to ESO. ACTION: NJC to send details of the suggestion to RGM. ACD, original: 10/12/16, revised: 15/12/16.